Toss a pair of dice. What are the odds against obtaining doubles? (i.e.,each die has the same number).
Answer by stanbon(75887)

Toss a pair of dice. What are the odds against obtaining doubles?
Prob(doubles)=6/36 = 1/6
Prob (not doubles) = 5/6
Odds against doubles = (5/6)/(1/6) = 5:1
Cheers,
Stan H.